# Beyond Discretization: Learning the Optimal Solution Path

### v2

@@ 3 SGD to Learn the Solution Path
Consider a family of functions $g(\cdot,z):\mathbb{R}^{d}\to\mathbb{R}$, $z\sim\mathbb{P}_{z}$, and $G(\cdot):=\mathbb{E}_{z\sim\mathbb{P}_{z}}\left[g(\cdot,z)\right]$. Let $G^{*}:=\min_{\boldsymbol{w}\in\mathbb{R}^{d}}G(\boldsymbol{w})$. Then, $g$ and $G$ are said to satisfy the relaxed weak growth condition with constants $\rho\geq 0$ and $\sigma\geq 0$, if for all $\boldsymbol{w}\in\mathbb{R}^{d}$,
$\mathbb{E}_{z\sim\mathbb{P}_{z}}\left[\|\nabla g(\boldsymbol{w},z)\|^{2}\right]~{}\leq~{}2\rho(G(\boldsymbol{w})-G^{*})+\sigma^{2}.$
@@
Suppose that the RWGC holds for the family of functions $g(\cdot,z):\mathbb{R}^{d}\to\mathbb{R}$, and that $G(\cdot):=\mathbb{E}_{z\sim\mathbb{P}_{z}}\left[g(\cdot,z)\right]$ is $\mu_{g}$-strongly convex for some for some $\mu_{g}>0$. Consider the SGD algorithm, initialized at $\boldsymbol{w}_{0}\in\mathbb{R}^{d}$, with iterations
$\boldsymbol{w}_{t+1}=\boldsymbol{w}_{t}-\frac{\bar{\eta}}{\rho}\nabla g(\boldsymbol{w}_{t},z_{t}),$
where $z_{t}\sim\mathbb{P}_{z}$ is a random sample and $\bar{\eta}\in\left(0,\min\left\{1,\frac{\rho}{\mu_{g}}\right\}\right]$ parametrizes the step-size. Then for all $t\geq 0$, we have
$\mathbb{E}[\|\boldsymbol{w}_{t}-\boldsymbol{w}^{*}\|^{2}]~{}\leq~{}\left(1-\frac{\bar{\eta}\mu_{g}}{\rho}\right)^{t}\|\boldsymbol{w}_{0}-\boldsymbol{w}^{*}\|^{2}+\frac{\bar{\eta}\sigma^{2}}{\rho\mu_{g}}.$
@@ 3.1 Solving Equation 2 with SGD
$f(\boldsymbol{\beta},\lambda)$ $~{}:=~{}h(\boldsymbol{\Phi}(\lambda)\boldsymbol{\beta},\lambda),$ (4a)
$F(\boldsymbol{\beta})$ $~{}:=~{}\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[f(\boldsymbol{\beta},\tilde{\lambda})\right]$
$~{}=~{}\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{\beta},\tilde{\lambda})\right],$ (4b)
@@
Suppose that Assumptions 2.2 and 2.1 hold and recall the constants defined in Equation 3. Then, the family of functions $f(\cdot,\lambda)$ and the function $F(\cdot)$, defined in Equation 4, satisfy the relaxed weak growth condition (RWGC) with constants $\rho=CL$ and $\sigma^{2}=2CL\epsilon_{\rm sp}^{\ast}$. Namely, for all $\boldsymbol{\beta}\in\mathbb{R}^{d}$,
$\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\|\nabla f(\boldsymbol{\beta},\tilde{\lambda})\|^{2}\right]~{}\leq~{}2CL(F(\boldsymbol{\beta})-F^{*})+2CL\epsilon_{\rm sp}^{\ast}.$
@@
$\boldsymbol{\beta}_{t+1}\leftarrow\boldsymbol{\beta}_{t}-\eta\nabla_{\boldsymbol{\beta}}h(\boldsymbol{\Phi}(\tilde{\lambda}_{t})\boldsymbol{\beta}_{t},\tilde{\lambda}_{t}),$
@@
$\nabla_{\boldsymbol{\beta}}h(\boldsymbol{\Phi}(\tilde{\lambda}_{t})\boldsymbol{\beta}_{t},\tilde{\lambda}_{t})=\boldsymbol{\Phi}(\tilde{\lambda}_{t})^{\top}\nabla_{\boldsymbol{\theta}}h(\boldsymbol{\Phi}(\tilde{\lambda}_{t})\boldsymbol{\beta}_{t},\tilde{\lambda}_{t}),$
and $\tilde{\lambda}_{t}\sim\mathbb{P}_{\lambda}$. We assume that the gradient $\nabla_{\boldsymbol{\theta}}h(\boldsymbol{\theta},\lambda)$ is easily computable for any $\boldsymbol{\theta}$ and $\lambda$.
@@
$\mathbb{E}\left[\epsilon_{\rm sp}(\boldsymbol{\beta}_{T})\right]~{}\leq~{}$ $2CL\|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}\left(1-\frac{\bar{\eta}c\mu}{CL}\right)^{T}$
@@
In particular, when $\epsilon_{\rm sp}^{\ast}>0$, then for $T\geq\left\lceil\frac{CL}{\bar{\eta}c\mu}\log\left(\frac{c\mu\|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}}{2\bar{\eta}\epsilon_{\rm sp}^{\ast}}\right)\right\rceil$,
$\mathbb{E}\left[\epsilon_{\rm sp}(\boldsymbol{\beta}_{T})\right]~{}\leq~{}\left(\frac{8CL(\bar{\eta}+1)}{c\mu}\right)\epsilon_{\rm sp}^{\ast}.$
When $\epsilon_{\rm sp}^{\ast}=0$, then for any $\epsilon>0$ and $T\geq\left\lceil\frac{CL}{\bar{\eta}c\mu}\log\left(\frac{2CL\|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}}{\epsilon}\right)\right\rceil$,
$\mathbb{E}\left[\epsilon_{\rm sp}(\boldsymbol{\beta}_{T})\right]~{}\leq~{}\epsilon.$
Theorem 3.4 highlights the role of the basis functions in our results. First we see that as $T\rightarrow\infty$, the expected solution path error plateaus at a constant proportional to $\frac{CL}{c\mu}\epsilon^{\ast}_{\rm sp}$. As mentioned, $\epsilon^{\ast}_{\rm sp}$ measures the expressiveness of the basis, and we expect $\epsilon_{\rm sp}^{*}$ to be small as we add more basis functions. We interpret $\frac{CL}{c\mu}$ as a condition number for Equation 2, which also depends on the choice of basis through Equation 3. This constant increases as we grow the basis. Finaly, the iteration complexity scales linearly with this condition number. Thus, an “ideal” basis must navigate this tradeoff between $\epsilon^{\ast}_{\rm sp}$ and $C/c$.
@@ 4 Specialized Results for $\Lambda=[-1,1]$
We use a simple basis: we approximate each component $i=1,\ldots,d$ of $\theta^{*}_{i}(\cdot)$ by the first $q$ Legendre polynomials. Hence the total number of basis functions is $p=qd$. Recall, the Legendre polynomials form an orthogonal basis on $[-1,1]$ with respect to the uniform distribution, i.e., $\mathbb{E}_{\tilde{\lambda}\sim\text{Unif}[-1,1]}\left[P_{n}(\tilde{\lambda})P_{m}(\tilde{\lambda})\right]=\frac{2}{2n+1}\mathbb{I}\left\{n=m\right\}$, where $P_{n}$ and $P_{m}$ are the $n^{\text{th}}$ and $m^{\text{th}}$ Legendre polynomial, respectively. Since we approximate each component separately, the matrix $\boldsymbol{\Phi}(\lambda)\in\mathbb{R}^{d\times qd}$ is block-diagonal, with d blocks of size $1\times q$.
@@
Bounding $\epsilon^{*}_{\rm sp}$ depends on the properties $\boldsymbol{\theta}^{*}(\lambda)$. As a first example,
@@
Let $\Lambda=[-1,1]$. Suppose Assumption 2.1 holds and that there exists an integer $\nu\geq 0$ and constant $V>0$ such that for all $i=1,\ldots,d$, $\theta^{*}_{i}(\lambda)$ has $\nu$ derivatives and the $\nu^{\text{th}}$ derivative $\boldsymbol{\theta}_{i}^{\ast(\nu)}$ is $V$-Lipschitz continuous. Then, for any $q\geq\nu+1$, for the basis described above, $\epsilon^{*}_{\rm sp}\ \leq\ \frac{dL}{2}\left(\frac{2V}{\pi\nu(q-\nu)^{\nu}}\right)^{2}.$
The proof of Lemma 4.2 is constructive; we exhibit a polynomial with the given solution path error.
@@
Let $\Lambda=[-1,1]$, $\mathbb{P}_{\lambda}$ be the uniform distribution on $[-1,1]$. Then, under the conditions of Lemma 4.2, if we use $q=O(\epsilon^{\frac{1}{2(1-\nu)}})$ polynomials in the previous basis, and run constant-step size SGD for $O(\epsilon^{\frac{1}{1-\nu}}\log(1/\epsilon))$ iterations, the resulting iterate $\boldsymbol{\beta}_{T}$ satisfies $\mathbb{E}\left[\epsilon_{\rm sp}(\boldsymbol{\beta}_{T})\right]\leq\epsilon$.
For clarity, both big “Oh” terms should be interpreted as $\epsilon\rightarrow 0$, and both suppress constants not depending on $\epsilon$ (but possibly depending on $h$). Loosely, the theorem establishes that the “smoother” $\boldsymbol{\theta}^{*}(\cdot)$ is (i.e. larger $\nu$), the fewer iterations required by our method to achieve a target tolerance.
@@
This gap is particularly striking $\nu\rightarrow\infty$. As a second example, suppose $\theta^{*}_{i}(\lambda)$ is analytic on $[-1,1]$, i.e, its Taylor Series is absolutely convergent on this interval.
@@
Suppose Assumption 2.1 holds and that for each $i=1,\ldots,d$, $\theta^{*}_{i}(\cdot)$ is analytic on the interval $[-1,1]$. Then, there exist constants $\omega>1$ and $M>0$ such that for any $q>0$, with the basis described above, $\epsilon^{*}_{\rm sp}\leq\frac{dL}{2}\left(\frac{2M\omega^{-q}}{\omega-1}\right)^{2}.$
@@
Let $\Lambda=[-1,1]$, and $\mathbb{P}_{\lambda}$ be the uniform distribution on $[-1,1]$. Then, under the conditions of Lemma 4.4, if we use $q=\log(1/\epsilon)/\log(\omega)$ polynomials in the previous basis, and run constant-step size SGD for $O\left(\log^{2}(1/\epsilon)\log\log(1/\epsilon)\right)$ iterations, the resulting iterate $\boldsymbol{\beta}_{T}$ satisfies $\mathbb{E}\left[\epsilon_{\rm sp}(\boldsymbol{\beta}_{T})\right]\leq\epsilon$.
@@
Again, the big “Oh” hides constants that do not depend on $\epsilon$. Compared to Ndiaye et al. (2019), the amount of work almost exponentially smaller.
5 An Adaptive Algorithm
@@
Although the previous section’s specialized results provide theoretical insight, they do not suggest practical algorithms. Picking the number of basis apriori from approximation theory is difficult. This challenge motivates our algorithm Algorithm 1, where the number of basis functions is determined “on the fly.” To implement Algorithm 1, one needs to specify i) a sequence of basis functions ii) a distribution $\mathbb{P}_{\lambda}$ and iii) a criterion for deciding when to add a new basis function. We next discuss these choices.
@@
Although any basis and distribution can be used in Algorithm 1, we suggest making these choices in concert. In our experiments we focus on sequences of polynomials that are orthogonal with respect to $\mathbb{P}_{\lambda}$. The Legendre polynomials and the uniform distribution on $[-1,1]$ is one such pair, but there are many canonical examples including Hermite polynomials with the normal distribution and Laguerre polynomials with the exponential distribution. There are performance-optimized implementations of these families in standard software (see, e.g., scipy.special). Although polynomials map to $\mathbb{R}$, by approximating each dimension separately as in Section 4, we can extend these polynomials to a basis for $d>1$. Finally, polynomials are uniform function approximators for Lipschitz functions. Thus, the class is highly expressive.
Two other advantages of orthogonal polynomials in Algorithm 1 are: i) Extending the basis does not require altering existing basis functions. This is unlike, e.g., cubic splines or adding nodes to a neural network, where the addition of a new spline point affects previous bases. ii) Intuitively, orthogonality suggests that if $\hat{\boldsymbol{\beta}}_{p-1}$ is near optimal in the $(p-1)^{\text{th}}$ iteration, then $[\hat{\boldsymbol{\beta}}_{p-1},0]$ is likely to be close to optimal in the $p^{\text{th}}$ iteration. Thus, we benefit from warmstarts.
We next discuss the criterion for adding a new basis function. One approach is to empirically approximate the objective of Equation 2 using a hold-out validation set, and add a basis function when performance stalls. This approach requires additional functional evaluations and samples $\tilde{\lambda}$. In our experiments, we explore an alternate criterion motivated by the RWGC. Specifically, if $\boldsymbol{\beta}$ is near-optimal, the upperbound on the second moment of the gradient in RWGC is approximately constant. Thus, we add a basis function when second moment of the gradient averaged over recent iterations has plateaued. This allows us to reuse gradients and additional function evaluations.

### v3

@@ 3 SGD TO LEARN THE SOLUTION PATH
Consider a family of functions $g(\cdot,z):\mathbb{R}^{d}\to\mathbb{R}$, $z\sim\mathbb{P}_{z}$, and $G(\cdot):=\mathbb{E}_{z\sim\mathbb{P}_{z}}\left[g(\cdot,z)\right]$. Let $G^{*}:=\min_{\bm{w}\in\mathbb{R}^{d}}G(\bm{w})$. Then, $g$ and $G$ are said to satisfy the relaxed weak growth condition with constants $\rho\geq 0$ and $\sigma\geq 0$, if for all $\bm{w}\in\mathbb{R}^{d}$,
$\mathbb{E}_{z\sim\mathbb{P}_{z}}\left[\|\nabla g(\bm{w},z)\|^{2}\right]~{}\leq~{}2\rho(G(\bm{w})-G^{*})+\sigma^{2}.$
@@
Suppose that the RWGC holds for the family of functions $g(\cdot,z):\mathbb{R}^{d}\to\mathbb{R}$, and that $G(\cdot):=\mathbb{E}_{z\sim\mathbb{P}_{z}}\left[g(\cdot,z)\right]$ is $\mu_{g}$-strongly convex for some $\mu_{g}>0$. Consider the SGD algorithm, initialized at $\bm{w}_{0}\in\mathbb{R}^{d}$, with iterations
$\bm{w}_{t+1}=\bm{w}_{t}-\frac{\bar{\eta}}{\rho}\nabla g(\bm{w}_{t},z_{t}),$
where $z_{t}\sim\mathbb{P}_{z}$ is a random sample and $\bar{\eta}\in\left(0,\min\left\{1,\frac{\rho}{\mu_{g}}\right\}\right]$ parametrizes the step-size. Let $\bm{w}^{*}:=\arg\min_{\bm{w}\in\mathbb{R}^{d}}G(\bm{w})$. Then for all $t\geq 0$, we have
$\mathbb{E}[\|\bm{w}_{t}-\bm{w}^{*}\|^{2}]~{}\leq~{}\left(1-\frac{\bar{\eta}\mu_{g}}{\rho}\right)^{t}\|\bm{w}_{0}-\bm{w}^{*}\|^{2}+\frac{\bar{\eta}\sigma^{2}}{\rho\mu_{g}}.$
@@ 3.1 Solving Equation 2 with SGD
$f(\bm{\beta},\lambda)$ $:=h(\bm{\Phi}(\lambda)\bm{\beta},\lambda),$ (4a)
$F(\bm{\beta})$ $:=\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[f(\bm{\beta},\tilde{\lambda})\right]=\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\bm{\Phi}(\tilde{\lambda})\bm{\beta},\tilde{\lambda})\right],$
@@
Suppose that Assumptions 2.2 and 2.1 hold and recall the constants defined in Equation 3. Then, the family of functions $f(\cdot,\lambda)$ and the function $F(\cdot)$, defined in Equation 4, satisfy the relaxed weak growth condition (RWGC) with constants $\rho=CL$ and $\sigma^{2}=2CL\epsilon_{\rm sp}^{\ast}$. Namely, for all $\bm{\beta}\in\mathbb{R}^{d}$,
$\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\|\nabla f(\bm{\beta},\tilde{\lambda})\|^{2}\right]~{}\leq~{}2CL(F(\bm{\beta})-F^{*})+2CL\epsilon_{\rm sp}^{\ast}.$
@@
$\bm{\beta}_{t+1}\leftarrow\bm{\beta}_{t}-\eta\nabla_{\bm{\beta}}h(\bm{\Phi}(\tilde{\lambda}_{t})\bm{\beta}_{t},\tilde{\lambda}_{t}),$
@@
$\nabla_{\bm{\beta}}h(\bm{\Phi}(\tilde{\lambda}_{t})\bm{\beta}_{t},\tilde{\lambda}_{t})=\bm{\Phi}(\tilde{\lambda}_{t})^{\top}\nabla_{\bm{\theta}}h(\bm{\Phi}(\tilde{\lambda}_{t})\bm{\beta}_{t},\tilde{\lambda}_{t}),$
and $\tilde{\lambda}_{t}\sim\mathbb{P}_{\lambda}$. We assume that the gradient $\nabla_{\bm{\theta}}h(\bm{\theta},\lambda)$ is easily computable for any $\bm{\theta}$ and $\lambda$.
@@
$\mathbb{E}\left[\epsilon_{\rm sp}(\bm{\beta}_{T})\right]~{}\leq~{}$ $2CL\|\bm{\beta}_{0}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}\left(1-\frac{\bar{\eta}c\mu}{CL}\right)^{T}$
@@
In particular, when $\epsilon_{\rm sp}^{\ast}>0$, then for $T\geq\left\lceil\frac{CL}{\bar{\eta}c\mu}\log\left(\frac{c\mu\|\bm{\beta}_{0}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}}{2\bar{\eta}\epsilon_{\rm sp}^{\ast}}\right)\right\rceil$,
$\mathbb{E}\left[\epsilon_{\rm sp}(\bm{\beta}_{T})\right]~{}\leq~{}\left(\frac{8CL(\bar{\eta}+1)}{c\mu}\right)\epsilon_{\rm sp}^{\ast}.$
When $\epsilon_{\rm sp}^{\ast}=0$, then for any $\epsilon>0$ and $T\geq\left\lceil\frac{CL}{\bar{\eta}c\mu}\log\left(\frac{2CL\|\bm{\beta}_{0}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}}{\epsilon}\right)\right\rceil$,
$\mathbb{E}\left[\epsilon_{\rm sp}(\bm{\beta}_{T})\right]~{}\leq~{}\epsilon.$
Theorem 3.4 highlights the role of the basis functions in our results. First we see that as $T\rightarrow\infty$, the expected solution path error plateaus at a constant proportional to $\frac{CL}{c\mu}\epsilon^{\ast}_{\rm sp}$. As mentioned, $\epsilon^{\ast}_{\rm sp}$ measures the expressiveness of the basis, and we expect $\epsilon_{\rm sp}^{*}$ to be small as we add more basis functions. We interpret $\frac{CL}{c\mu}$ as a condition number for Equation 2, which also depends on the choice of basis through Equation 3. This constant increases as we grow the basis. Finally, the iteration complexity scales linearly with this condition number. Thus, an “ideal” basis must navigate this tradeoff between $\epsilon^{\ast}_{\rm sp}$ and $C/c$.
@@ 4 SPECIALIZED RESULTS FOR $\Lambda=[-1,1]$
We use a simple basis: we approximate each component $i=1,\ldots,d$ of $\theta^{*}_{i}(\cdot)$ by the first $q$ Legendre polynomials. Hence the total number of basis functions is $p=qd$. Recall, the Legendre polynomials form an orthogonal basis on $[-1,1]$ with respect to the uniform distribution, i.e., $\mathbb{E}_{\tilde{\lambda}\sim\text{Unif}[-1,1]}\left[P_{n}(\tilde{\lambda})P_{m}(\tilde{\lambda})\right]=\frac{2}{2n+1}\mathbb{I}\left\{n=m\right\}$, where $P_{n}$ and $P_{m}$ are the $n^{\text{th}}$ and $m^{\text{th}}$ Legendre polynomial, respectively. Since we approximate each component separately, the matrix $\bm{\Phi}(\lambda)\in\mathbb{R}^{d\times qd}$ is block-diagonal, with d blocks of size $1\times q$.
@@
Bounding $\epsilon^{*}_{\rm sp}$ depends on the properties $\bm{\theta}^{*}(\lambda)$. As a first example,
@@
Let $\Lambda=[-1,1]$. Suppose Assumption 2.1 holds. Further assume that there exists an integer $\nu\geq 0$ and constant $V>0$ such that for all $i=1,\ldots,d$, $\bm{\theta}^{*}_{i}(\cdot)$ has $\nu$ derivatives, where $\bm{\theta}^{*}_{i}(\lambda)$, …, $\bm{\theta}^{(\nu-1)*}_{i}(\lambda)$ are absolutely continuous and $\bm{\theta}^{(\nu)*}_{i}(\lambda)$ has total variation bounded by $V$. Then, for any $q\geq\nu+1$, for the basis described above, $\epsilon^{*}_{\rm sp}\ \leq\ \frac{dL}{2}\left(\frac{2V}{\pi\nu(q-\nu)^{\nu}}\right)^{2}.$
The proof of Lemma 4.2 is constructive; we exhibit a polynomial with the given solution path error. The bound confirms the intuition that if $\bm{\theta}^{*}(\cdot)$ is smooth (has many derivatives), then adding basis function drives down the approximation error rapidly. Problems arising in many application domains, including machine learning problems like ridge regression and relatives, often exhibit highly or even infinitely differentiable solution paths.
@@
Let $\Lambda=[-1,1]$, $\mathbb{P}_{\lambda}$ be the uniform distribution on $[-1,1]$. Then, under the conditions of Lemma 4.2 and assume $\nu\geq 2$, if we use $q=O(\epsilon^{\frac{1}{2(1-\nu)}})$ polynomials in the previous basis, and run constant-step size SGD for $O(\epsilon^{\frac{1}{1-\nu}}\log(1/\epsilon))$ iterations, the resulting iterate $\bm{\beta}_{T}$ satisfies $\mathbb{E}\left[\epsilon_{\rm sp}(\bm{\beta}_{T})\right]\leq\epsilon$.
For clarity, while Lemma 4.2 holds for any $\nu\geq 0$, Theorem 4.3 requires $\nu\geq 2$. Moreover, both big “Oh” terms should be interpreted as $\epsilon\rightarrow 0$, and both suppress constants not depending on $\epsilon$ (but possibly depending on $h$). The theorem establishes that the “smoother” $\bm{\theta}^{*}(\cdot)$ is (i.e. larger $\nu$), the fewer iterations required by our method to achieve a target tolerance.
@@
This gap becomes more striking as $\nu\rightarrow\infty$. Hence, we next consider the case where $\theta^{*}_{i}(\lambda)$ is analytic on $[-1,1]$, i.e, its Taylor Series is absolutely convergent on this interval.
@@
Suppose Assumption 2.1 holds and that for each $i=1,\ldots,d$, $\bm{\theta}^{*}_{i}(\cdot)$ is analytic on the interval $[-1,1]$. Then, there exist constants $\omega>1$ and $M>0$ such that for any $q>0$, with the basis described above, $\epsilon^{*}_{\rm sp}\leq\frac{dL}{2}\left(\frac{2M\omega^{-q}}{\omega-1}\right)^{2}.$
@@
Let $\Lambda=[-1,1]$, and $\mathbb{P}_{\lambda}$ be the uniform distribution on $[-1,1]$. Then, under the conditions of Lemma 4.4, if we use $q=\log(1/\epsilon)/\log(\omega)$ polynomials in the previous basis, and run constant-step size SGD for $O\left(\log^{2}(1/\epsilon)\log\log(1/\epsilon)\right)$ iterations, the resulting iterate $\bm{\beta}_{T}$ satisfies $\mathbb{E}\left[\epsilon_{\rm sp}(\bm{\beta}_{T})\right]\leq\epsilon$.
@@
Again, the big “Oh” hides constants that do not depend on $\epsilon$. Compared to Ndiaye et al. (2019), the amount of work required is almost exponentially smaller for $\dim(\Lambda)=1$.
In higher dimensions, the curse of dimensionality causes the required discretization points to grow exponentially with $\dim(\Lambda)$. In contrast, our method’s performance depends not on $\dim(\Lambda)$ but on the minimal solution path error $\epsilon_{\mathrm{sp}}^{*}$. The key advantage arises when we either select a basis ensuring a small $\epsilon_{\mathrm{sp}}^{*}$ using prior knowledge or adopt a highly flexible, adaptive approach, as we next discuss in Section 5.
5 IMPLEMENTATION GUIDELINES
@@
Although Section 4 provides insight on how to select a basis (in certain cases), choosing the “right” basis apriori from approximation theory remains a difficult challenge. To circumvent this issue, we re-introduce Algorithm 1. Algorithm 1 is more practical, as we only need to specify: i) a sequence of basis functions, ii) a distribution $\mathbb{P}_{\lambda}$ and, iii) a criterion for deciding when to add a new function to the basis. We next provide intuition and practical guidance on these choices.
Intuition for an “Ideal” Basis. Consider
$h(\bm{\theta},\bm{\lambda})\ =\ \bm{\theta}^{\top}\bm{Q}(\bm{\lambda})\bm{\theta}+\bm{b}(\bm{\lambda})^{\top}\bm{\theta}.$
In this special case, Equation 2 becomes
$\min_{\bm{\beta}}\bm{\beta}^{\top}\mathbb{E}\left[\bm{\Phi}(\bm{\tilde{\lambda}})^{\top}\bm{Q}(\bm{\tilde{\lambda}})\bm{\Phi}(\bm{\tilde{\lambda}})\right]\bm{\beta}+\mathbb{E}\left[\bm{b}(\bm{\tilde{\lambda}})^{\top}\bm{\Phi}(\bm{\tilde{\lambda}})\right]\bm{\beta}.$
The complexity of this problem depends strongly on its condition number (in our theory, this is bounded by $\frac{CL}{c\mu}$), which equals the condition number of the matrix $\mathbb{E}\left[\bm{\Phi}(\bm{\tilde{\lambda}})^{\top}\bm{Q}(\bm{\tilde{\lambda}})\bm{\Phi}(\bm{\tilde{\lambda}})\right]$ and is directly computable in this case. Hence, an ideal basis would ensure this condition number is as close to $1$ (its lower bound) as possible. In particular, if the basis functions $\bm{\Phi}_{j}(\cdot)$ are orthonormal with respect to the inner product
$\langle\bm{\Phi}_{j}(\cdot),\bm{\Phi}_{k}(\cdot)\rangle\ \equiv\ \mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[\bm{\Phi}_{j}(\bm{\tilde{\lambda}})^{\top}\bm{Q}(\bm{\tilde{\lambda}})\bm{\Phi}_{k}(\bm{\tilde{\lambda}})\right],$
then, the resulting condition number is $1$. For general $h$, insofar as $F(\bm{\beta})=\mathbb{E}_{\tilde{\lambda}\sim\mathbb{P}_{\lambda}}\left[h(\bm{\Phi}(\tilde{\lambda})\bm{\beta},\tilde{\lambda})\right]$ might be approximated by a second-order Taylor series expansion around the optimum $\bm{\beta}_{\mathrm{avg}}^{\ast}$, a good basis might be orthonormal w.r.t. to the Hessian of $F(\cdot)$ at optimality. We verify this in Section 6.3. Absent knowledge about the Hessian at optimality, the best approach is to use a family of flexible and expressive basis functions. In Sections 6.1 and 6.2, we use orthogonal polynomials for the following reasons: i) Adding new basis functions does not require altering the existing basis, and ii) Intuitively, orthogonality suggests that if $\hat{\bm{\beta}}_{p-1}$ is near optimal in the $(p-1)^{\text{th}}$ iteration, then $[\hat{\bm{\beta}}_{p-1},0]$ is likely to be a near optimal (and good warm-start) solution in the $p^{\text{th}}$ iteration. The first benefit is not shared, for example, by other approaches like cubic splines or neural networks.
@@
Interplay Between Basis and Distribution. Although any sequence of basis functions and any distribution can be used in Algorithm 1, we suggest making these choices in concert. In our experiments, we focus on sequences of polynomials that are orthogonal with respect to $\mathbb{P}_{\lambda}$. The Legendre polynomials and the uniform distribution on $[-1,1]$ is one such pair, but there are many canonical examples including Hermite polynomials with the normal distribution and Laguerre polynomials with the exponential distribution. There are performance-optimized implementations of these families in standard software (see, e.g., scipy.special). Although polynomials map to $\mathbb{R}$, by approximating each dimension separately as in Section 4, we can extend polynomials to a basis for $d>1$. Finally, polynomials are highly expressive as they are uniform approximators for Lipschitz functions.
Deciding the Number of Basis Functions “On the Fly”. As reflected by our theoretical convergence guarantees, choosing a large, flexible basis induces a tradeoff. A larger basis increases the condition number of Equation 2 (as measured by the ratio $\frac{CL}{c\mu}$), and hence convergence will be slower. This observation motivates our adaptive approach that only adds basis functions as necessary until we reach a desired accuracy. One approach is to empirically approximate the objective of Equation 2 using a hold-out validation set, and add a basis function when performance stalls. In our experiment with low dimensional hyperparameters (Section 6.1 and 6.2), we use Gauss-Legendre quadrature to evaluate the stochastic objective exactly over the entire hyperparameter space. In the high dimensional hyperparameter experiment (Section 6.3), we instead compute the performance of our learned $\bm{\beta}$ through an ERM version of the stochastic objective on a validation set of size $1000$.